OpenStudy (anonymous):

What are the steps when finding the distance between (-2, 4) and (7, -6)?

OpenStudy (anonymous):

find the linear distance between each of the two coordinates by subtraction then use pythagoras disance between \(-2\) and \(7\) is \(7+2=9\) distance between \(4\) and \(-6\) is \(4+6=10\) pythagoras gives tells you \(a^2+b^2=d^2\) so \[d^2=9^2+10^2\] and therefore \[d=\sqrt{9^2+10^2}=\sqrt{181}\]

OpenStudy (anonymous):

many people like formulas, it is \[d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]